40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 16 03:43:09 2014, Jenne, Update, LSC, PRMI loops need help PRMI_55vs165_15Dec2014.pdf
    Reply  Tue Dec 16 20:51:18 2014, diego, Update, LSC, PRMI loops need help 
    Reply  Wed Dec 17 13:14:38 2014, ericq, Update, LSC, PRMI loops need help oops.pdf
       Reply  Wed Dec 17 14:42:13 2014, Jenne, Update, LSC, PRMI loops need help 
Message ID: 10809     Entry time: Wed Dec 17 13:14:38 2014     In reply to: 10804     Reply to this: 10810
Author: ericq 
Type: Update 
Category: LSC 
Subject: PRMI loops need help 

Quote:

However, the PRMI would not acquire lock with the arms held off resonance. 

This is entirely my fault. 

Last week, while doing some stuff with PRY, I put this filter in SUS_PRM_LSC, to stop some saturations from high frequency sensing noise

oops.pdf

After the discussion at today's meeting, it struck me that I might have left it on. Turns out I did. 

20 degree phase lag at 200Hz can explain the instability, some non-flat shape at few hundreds of Hz explains the non 1/f shape.

Sorry about all that...

ELOG V3.1.3-